# Entanglement Spectrum Dynamics as a Probe for Non-Hermitian Bulk-Boundary Correspondence in Systems with Periodic Boundaries

### v1

@@ Abstract
It has recently been established that open quantum systems may exhibit a strong spectral sensitivity to boundary conditions, known as the non-Hermitian/Liouvillian skin effect (NHSE/LSE), making the topological properties of the system boundary-condition sensitive. In this Letter, we ask the query: Can topological phase transitions of open quantum systems, captured by open boundary conditioned invariants, be observed in the dynamics of a system in a periodic boundary condition, even in the presence of NHSE/LSE? We affirmatively respond to this question, by considering the quench dynamics of entanglement spectrum in a periodic open quantum fermionic system. We demonstrate that the entanglement spectrum exhibits zero-crossings only when this periodic system is quenched from a topologically trivial to non-trivial phase, defined from the spectrum in open boundary conditions, even in systems featuring LSE. Our results reveal that non-Hermitian topological phases leave a distinctive imprint on the unconditional dynamics within a subsystem of fermionic systems.
Introduction.- The recent expansion of phases of matter into the non-Hermitian realm has significantly broadened our understanding of free fermionic systems [1, 2, 3]. The presence of non-Hermiticity, which generally arises from the exchange of energy and/or particles of a system with the environment, leads to various exotic phenomena. A paradigmatic example is the emergence of extreme spectral sensitivity to boundary conditions, often referred to as the non-Hermitian skin effect (NHSE) [4, 6, 5, 11, 14, 10, 12, 13], or Liouvillian skin effect (LSE) in the context of open quantum systems [6, 7, 18, 19, 20, 21], see also [8, 9] for related results in interacting many-body systems. In the presence of NHSE/LSE, the system exhibits exponential localization of eigenstates near its boundaries. This poses a challenge to the conventional understanding of bulk-boundary correspondence: the presence of topological edge states in open boundary conditions (OBC) cannot be inferred from the topological invariant of the Bloch Hamiltonian [4, 5]. In this situation, one must extend the conventional Bloch band theory to incorporate complex-valued wave vectors in order to identify topological invariants in OBC [4, 5]. This is known as the non-Hermitian bulk-boundary correspondence.
@@
In this Letter, in sharp contrast to this naïve expectation, we show that there are cases where OBC topological invariants can be extracted from the dynamics of the PBC system. This is achieved by examining the quench dynamics of the entanglement spectrum (ES). Initially recognized in the context of a fractional quantum Hall insulator [24] and later formalized for free fermionic [25] and one-dimensional systems [26], the ES has since played a pivotal role in understanding topological phases of both non-interacting and interacting systems. In highly non-equilibrium states, the ES serves as a dynamical probe for wavefunction topology [27, 29, 28]. Recent investigations have even extended its applicability to open quantum systems, where the real line-gap topology was shown to be detectable through the quench dynamics of ES [30, 31, 32]. However, these works considered systems that do not exhibit LSE [30, 31, 32], leaving the effect of boundary conditions sensitivity on the dynamics of ES unclear.
We examine in this Letter an open quantum fermionic lattice exhibiting LSE and show that the dynamics of the ES can serve as a probe for topological phases defined from the OBC spectrum, even when the physical system is in PBC. In particular, we numerically analyze an open quantum model whose dynamics is governed by a dynamical matrix identical to the non-Hermitian Su-Schrieffer-Heeger (SSH) model, a paradigmatic model exhibiting NHSE/LSE [4, 6, 5, 19, 21, 21] and demonstrate that in PBC systems, topological phases defined from the OBC spectrum of Lindbladians are discernible through mode crossings of the ES during quench dynamics. As the ES in open free fermion systems can be computed from a two-point correlation function of a subsystem [36], our proposal to detect the non-Hermitian bulk-boundary correspondence does not require post-selection.
@@
Figure 2: (a) The system is initially prepared in the trivial ground state of the SSH model  [33] in PBC with no dissipation. Then, at $t>0$, the relevant system $(\mathrm{S}+\bar{\mathrm{S}})$ is driven out of its initial state by an abrupt change of the hopping parameters $J_{1},J_{2}$ and a correlated dissipation with rate $\gamma_{1}$ given by Eq. (8) is simultaneously turned on. The relevant system evolves according to the GKLS equation Eq. (1) in PBC. (b) Phase diagram of the Lindbladian SSH model Eq. (11). Topological phase transitions are separated by the solid and dotted lines for OBC and PBC, respectively. (c), (d) ES dynamics for a subsystem S. (c) Trivial to topological quenches for both OBC and PBC spectrum ($\omega_{\rm OBC}=1,\omega_{\rm PBC}=1/2$) $(J_{1}/J_{2},\gamma_{1}/(2J_{2}))=(0.7,0.5)$. The green lines represent the four eigenvalues, $\xi_{l}(t)$, of the covariance matrix $\mathsf{C}_{\rm SS}(t)$ that are closest to the zero energy entanglement value $\xi=0.5$ during the time evolution. The other eigenvalues are shown in blue. Four ES eigenvalues cross $\xi$ = 0.5 as they decay. (d) Trivial to trivial quenches for OBC spectrum ($\omega_{\rm OBC}=0$$(J_{1}/J_{2},\gamma_{1}/(2J_{2}))=(1.5,0.8)$ but topological for PBC spectrum ($\omega_{\rm PBC}=1/2$). All the eigenvalues, $\xi_{l}(t)$, of $\mathsf{C}_{\rm SS}(t)$ are shown in blue. All eigenvalues decay to $\xi=0.5$ without crossing it. For both cases, we set the initial state to be the (trivial) ground state at $(J_{1}/J_{2},\gamma_{1}/(2J_{2}))=(1.5,0)$, the system size $L=20$, and the size of subsystem S is $M=10$ sites. Note crucially that, while the ES dynamics reflect the OBC spectrum, our physical system is in PBC.
@@
The model.- We investigate the dynamics of an open quantum system, illustrated in Fig. 1, where a free fermionic system ($\mathrm{S}+\bar{\mathrm{S}}$) is linearly coupled to a Markovian environment (${\rm env}$). After integrating out the environmental degrees of freedom, the reduced density matrix $\hat{\rho}={\rm Tr}_{\rm env}\hat{\rho}_{\rm tot}$ is governed by a quadratic Gorini–Kossakowski–Sudarshan–Lindblad(GKSL) equation [22, 23], given by,
$\partial_{t}\hat{\rho}(t)$ $=-i[\hat{H},\hat{\rho}(t)]+\sum_{\alpha}\mathcal{D}[\hat{L}_{\alpha}]\hat{\rho}(t)+\sum_{\alpha}\mathcal{D}[\hat{G}_{\alpha}]\hat{\rho}(t),$ (1)
@@
where $\hat{H}=\sum_{i,j=1}^{L}H_{i,j}\hat{c}_{i}^{\dagger}\hat{c}_{j}$ is the Hamiltonian of the relevant system $(\textrm{S}+\bar{\textrm{S}})$. Here, $\hat{c}_{i}^{\dagger}$ and $\hat{c}_{i}$ are creation and annihilation operators of fermions at site $i$, respectively. The dissipation arising from the coupling of the system to the Markovian reservoir is described by $\mathcal{D}[\hat{M}]\hat{\rho}:=\hat{M}\hat{\rho}\hat{M}^{\dagger}-\frac{1}{2}\{\hat{M}^{\dagger}\hat{M},\hat{\rho}(t)\}$. Each coupling to the Markovian reservoir, labeled by $\alpha$ (see Fig. 2(a)), induces incoherent addition or removal of particles and is described by linear jump operators, $\hat{L}_{\alpha}=\sum_{i}\ell_{i}^{\alpha}\hat{c}_{i}$ and $\hat{G}_{\alpha}=\sum_{i}g^{*\alpha}_{i}\hat{c}^{\dagger}_{i}$, respectively, where $\ell_{n}^{\alpha}$ and $g^{\alpha}_{n}$ are complex numbers.
@@
ES crossings as a non-Hermitian bulk-boundary correspondence indicator.— To dynamically detect the non-Hermitian bulk-boundary correspondence, we propose examining the ES dynamics after a dissipative quench. As depicted in Fig. 1, we partition the relevant system $(\mathrm{S}+\bar{\mathrm{S}})$ further into the subsystem ${\mathrm{S}}=\{1,2,...,M\}$ and its complement $\bar{\mathrm{S}}=\{M+1,M+2,...,L\}$. The reduced density matrix of the subsystem can be written as $\hat{\rho}_{\mathrm{S}}={\rm Tr}_{\bar{\mathrm{S}}}\hat{\rho}=(1/Z_{\mathrm{S}})e^{-\hat{H}_{\rm S}}$ with $Z_{\mathrm{S}}={\rm Tr}_{\mathrm{S}}[e^{-\hat{H}_{\rm S}}]$, where ${\rm Tr}_{\mathrm{S}(\bar{\mathrm{S}})}$ denotes the trace over the subregion S(${\mathrm{S}}(\bar{\mathrm{S}})$. We refer to the spectrum of $\hat{H}_{\rm S}$ as the ES $\epsilon_{l}$, the central quantity of interest. Using the property that $\hat{\rho}_{\rm S}$ is Gaussian, the entanglement Hamiltonian can be decomposed as $\hat{H}_{\rm S}=\sum_{l}\epsilon_{l}\hat{f}_{l}^{\dagger}\hat{f}_{l}$, where $\hat{f}_{l}$ is a fermionic annihilation operator of the $l$-th eigenmode. In the absence of the system-environment coupling, this recovers the conventional definition of the ES for closed fermionic systems, which is known to successfully detect bulk-boundary correspondence [24, 25, 26].
@@
References [27, 29, 28, 30, 31, 32] considered the situation where they first prepared a state corresponding to the ground state of a Hamiltonian in the trivial phase followed by a sudden change in the parameters. They showed that when the post-quench Hamiltonian has a topologically non-trivial (trivial) ground state, (no) zero-crossings of the ES occur.
@@
We investigate below a similar setup where we also quench the system from a topologically trivial to a non-trivial or trivial state in an open quantum system governed by Eq. (1). Importantly, the topology of our system is characterized by the non-Hermitian effective Hamiltonian $\hat{H}_{\rm eff}$, given by Eq. (2), that exhibits NHSE. We find that an analogous phenomenon to the closed-system counterpart occurs, which surprisingly reflects the topology of an OBC system even when the system itself is in PBC: physically traversing the system is unnecessary to detect topological phases of $\hat{H}_{\rm eff}$ defined under open boundary conditions.
@@
Practically, in our quadratic system, the ES can be computed from the covariance matrix $\mathsf{C}(t)$, whose $(j,i)$ component is given by the spatial correlation function $[{\sf C}]_{j,i}(t):={\rm Tr}_{\mathrm{S}+\bar{\mathrm{S}}}[\hat{\rho}(t)\hat{c}^{\dagger}_{i}\hat{c}_{j}]$ [36]. By taking advantage of the property that the density matrix is always Gaussian (as long as we take the initial state to be Gaussian) and Wick’s theorem applies as a result, the ES can be derived from the spectrum of the covariance matrix ${\sf C}_{\sf SS}$ in the subsystem S defined as the correlation matrix $\sf C$ whose indices are restricted to the region S,
$[{\sf C}_{\sf SS}]_{j,i}(t):={\rm Tr}_{\mathrm{S}+\bar{\mathrm{S}}}[\hat{\rho}(t)\hat{c}^{\dagger}_{i}\hat{c}_{j}];\;(i,j)\in{\sf S},$ (3)
@@
which can be expressed using the reduced density matrix of the subsystem S as $[{\sf C}_{\sf SS}]_{j,i}(t)={\rm Tr}_{\rm S}[\hat{\rho}_{\rm S}(t)\hat{c}^{\dagger}_{i}\hat{c}_{j}]$. Letting $\xi_{l}(t)$ be the eigenspectrum of $\mathsf{C}_{\rm SS}(t)$, one finds the relation between $\xi_{l}(t)$ and the ES $\epsilon_{l}(t)$  [36] (See Supplemental Materials (SM) [35]):
@@
Lindbladian SSH model.- To be specific, we consider the following system exhibiting LSE [4, 21] (Fig. 2(a)). For the system Hamiltonian, we consider the tight-binding Hamiltonian
$\hat{H}=\sum\limits_{j}\left(J_{1}\hat{c}_{A,j}^{\dagger}\hat{c}_{B,j}+J_{2}\hat{c}^{\dagger}_{B,j}\hat{c}_{A,j+1}+J_{3}\hat{c}_{A,j+1}^{\dagger}\hat{c}_{B,j}\right)+h.c.$ (6)
For $J_{3}=0$, the Hamiltonian reduces to the SSH model [33]. The dissipation of our model is given by the following jump operators for $j=1,2,\cdots,N$ (see Fig. 2 (a)):
$\hat{L}_{j,1}=\sqrt{\frac{\gamma_{1}}{2}}\left(\hat{c}_{A,j}+i\hat{c}_{B,j}\right),$ (7)
$\hat{G}_{j,1}=\sqrt{\frac{\gamma_{1}}{2}}(\hat{c}_{A,j}^{\dagger}-i\hat{c}_{B,j}^{\dagger}),$ (8)
$\hat{L}_{j,2}=\sqrt{\frac{\gamma_{2}}{2}}\left(\hat{c}_{A,j+1}+i\hat{c}_{B,j}\right),$ (9)
$\hat{G}_{j,2}=\sqrt{\frac{\gamma_{2}}{2}}(\hat{c}_{A,j+1}^{\dagger}-i\hat{c}_{B,j}^{\dagger}).$ (10)
Here, we have chosen the dissipators such that $\hat{L}_{j,s=1,2}^{\dagger}=\hat{G}_{j,s=1,2}$, ensuring the system reaches an infinite-temperature state at the long-time limit. The ES asymptotically converges to zero $\epsilon_{n}=0(\xi_{n}=0.5)$, avoiding the ES from crossing zero for trivial reasons with no topological origin [30, 37].
Equations (6)-(10) yield the effective Hamiltonian
@@
$+J_{3}\hat{c}_{A,j+1}^{\dagger}\hat{c}_{B,j}\Big{)}+h.c.\Big{]}+\sum_{j}\Big{(}\frac{\gamma_{1}}{2}(\hat{c}_{A,j}^{\dagger}\hat{c}_{B,j}-\hat{c}^{\dagger}_{B,j}\hat{c}_{A,j})$
$+\frac{\gamma_{2}}{2}(\hat{c}^{\dagger}_{B,j}\hat{c}_{A,j+1}-\hat{c}_{A,j+1}^{\dagger}\hat{c}_{B,j})\Big{)}-i\sum_{i}\gamma_{i}\hat{n},$ (11)
with $\hat{n}:=\sum_{j}[\hat{c}^{\dagger}_{A,j}\hat{c}_{A,j}+\hat{c}^{\dagger}_{B,j}\hat{c}_{B,j}]$, which is the non-Hermitian SSH model well studied in the literature (except for the constant shift in the imaginary axis that does not affect its topological features) that is known to exhibit NHSE [33]. For PBC, the Hamiltonian $\hat{H}_{\rm eff}$ can expressed as $\hat{H}_{\rm eff}=\sum_{k}\bm{c}^{\dagger}(k)\mathsf{H}_{\rm eff}(k)\bm{c}(k)$, where $\bm{c}(k):=(\hat{c}_{A}(k),\hat{c}_{B}(k))^{T}$ with $\hat{c}_{A(B)}(k)=1/\sqrt{L}\sum_{j}e^{-ikj}\hat{c}_{A(B),j}$ and $\mathsf{H}_{\rm eff}(k)=h_{1}(k)\sigma_{x}+h_{2}(k)\sigma_{y}$, where $h_{1}(k)=J_{1}+(J_{2}+J_{3})\cos(k)-i\gamma_{2}/2\sin(k)$, $h_{2}(k)=(J_{2}-J_{3})\sin(k)+i\gamma_{2}/2\cos(k)-i\gamma_{1}/2$. Here, $\sigma_{i}$ with $i\in\{x,y,z\}$ are the Pauli matrices. Since the effective Hamiltonian has a sublattice symmetry $\sigma_{z}\mathsf{H}_{\rm eff}(k)\sigma_{z}^{-1}=-\mathsf{H}_{\rm eff}(k)$, we can define the ${\mathbb{Z}}$ topological number of $\mathsf{H}_{\rm eff}(k)$, given by the winding number $\omega_{\rm PBC}=\oint\frac{dk}{4\pi i}{\rm Tr}[\sigma_{z}\mathsf{H}_{\rm eff}^{-1}(k)\frac{d}{dk}\mathsf{H}_{\rm eff}(k)]$ [2]. The winding number can be similarly defined for OBC, $\omega_{\rm OBC}$, by generalizing the momentum integral in the PBC winding number, $\omega_{\rm PBC}$, to those with complex values along the generalized Brillouin zone. (See e.g., Refs. [5, 14] for details.)
@@
In our simulation, we set the pre-quench state to be the ground state in the trivial phase of Eq. (11), in the Hermitian limit $\gamma_{1}=0$.

### v2

@@ Abstract
It has recently been established that open quantum systems may exhibit a strong spectral sensitivity to boundary conditions, known as the non-Hermitian/Liouvillian skin effect (NHSE/LSE), making the topological properties of the system boundary-condition sensitive. In this Letter, we ask the query: Can topological phase transitions of open quantum systems, captured by open boundary conditioned invariants, be observed in the dynamics of a system under periodic boundary conditions, even in the presence of NHSE/LSE? We affirmatively respond to this question, by considering the quench dynamics of entanglement spectrum in a periodic open quantum fermionic system. We demonstrate that the entanglement spectrum exhibits zero-crossings only when this periodic system is quenched from a topologically trivial to non-trivial phase, defined from the spectrum under open boundary conditions, even in systems featuring LSE. Our results reveal that non-Hermitian topological phases leave a distinctive imprint on the unconditional dynamics within a subsystem of fermionic systems.
Introduction.— The recent expansion of phases of matter into the non-Hermitian realm has significantly broadened our understanding of free fermionic systems [1, 2, 3]. The presence of non-Hermiticity, which generally arises from the exchange of energy and/or particles of a system with the environment, leads to various exotic phenomena. A paradigmatic example is the emergence of extreme spectral sensitivity to boundary conditions, often referred to as the non-Hermitian skin effect (NHSE) [4, 5, 6, 7, 8, 13, 14, 1, 2, 12, 11, 17, 20, 15, 16, 18, 19, 21, 22, 23, 3], or Liouvillian skin effect (LSE) in the context of open quantum systems [7, 8, 1, 2, 12, 11], see also [13, 14] for related results in interacting many-body systems. In the presence of NHSE/LSE, the system exhibits exponential localization of eigenstates near its boundaries. This poses a challenge to the conventional understanding of bulk-boundary correspondence: the presence of topological edge states in open boundary conditions (OBC) cannot be inferred from the topological invariant of the Bloch Hamiltonian [5, 6]. In this situation, one must extend the conventional Bloch band theory to incorporate complex-valued wave vectors in order to identify topological invariants in OBC [5, 6]. This is known as the non-Hermitian bulk-boundary correspondence.
@@
In this Letter, in sharp contrast to this naïve expectation, we show that there are cases where OBC topological invariants can be extracted from the dynamics of the PBC system. This is achieved by examining the quench dynamics of the entanglement spectrum (ES). Initially recognized in the context of a fractional quantum Hall insulator [27] and later formalized for free fermionic [28] and one-dimensional interacting systems [29], the ES has since played a pivotal role in understanding topological phases of both non-interacting and interacting systems. In highly non-equilibrium states, the ES serves as a dynamical probe for wavefunction topology [30, 32, 31]. Recent investigations have partially extended its applicability to non-Hermitian and open quantum systems. In non-Hermitian systems, which effectively describe the dynamics of open systems subject to post-selection, it was reported that for point-gapped systems, which exhibit NHSE, ES fails to capture topological properties of the ground state [33]. In open quantum systems, the real line-gap topology was shown to be detectable through the quench dynamics of ES [4, 35, 36]. However, these works on open quantum systems, considered systems that do not exhibit LSE [4, 35, 36], leaving the effect of boundary conditions sensitivity on the dynamics of ES unclear.
We examine in this Letter an open quantum fermionic lattice exhibiting LSE and show that the dynamics of the ES can serve as a probe for topological phases defined from the OBC spectrum, even when the physical system is in PBC. In particular, we numerically analyze an open quantum model whose dynamics is governed by a dynamical matrix identical to the non-Hermitian Su-Schrieffer-Heeger (SSH) model, a paradigmatic model exhibiting NHSE/LSE [5, 7, 6, 2, 11, 11] and demonstrate that in PBC systems, topological phases defined from the OBC spectrum of Lindbladians are discernible through mode crossings of the ES during quench dynamics. As the ES in open free fermionic systems can be computed from a two-point correlation function of a subsystem [40], our proposal to detect the non-Hermitian bulk-boundary correspondence does not require post-selection.
@@
Figure 2: (a) The system is initially prepared in the trivial ground state of the SSH model  [37] under PBC with no dissipation. Then, at $t>0$, the relevant system $(\mathrm{S}+\bar{\mathrm{S}})$ is driven out of its initial state by an abrupt change of the hopping parameters $J_{1},J_{2}$ and a correlated dissipation with rate $\gamma_{1}$ is simultaneously turned on. The relevant system evolves according to the GKLS equation Eq. (1) under PBC. (b) Phase diagram of the Lindbladian SSH model Eq. (7). Topological phase transitions are separated by the solid and dotted lines for OBC and PBC, respectively. (c), (d) ES dynamics for a subsystem S. (c) Trivial to topological quenches for both OBC and PBC spectrum ($\omega_{\rm OBC}=1,\omega_{\rm PBC}=1/2$) $(J_{1}/J_{2},\gamma_{1}/(2J_{2}))=(0.7,0.5)$. The green lines represent the four eigenvalues, $\xi_{l}(t)$, of the covariance matrix $\mathsf{C}_{\rm SS}(t)$ that are closest to the zero energy entanglement value $\xi=0.5$ during the time evolution. The other eigenvalues are shown in blue. Four ES eigenvalues cross $\xi$ = 0.5 as they decay. (d) Trivial to trivial quenches for OBC spectrum ($\omega_{\rm OBC}=0$$(J_{1}/J_{2},\gamma_{1}/(2J_{2}))=(1.5,0.8)$ but topological for PBC spectrum ($\omega_{\rm PBC}=1/2$). All the eigenvalues, $\xi_{l}(t)$, of $\mathsf{C}_{\rm SS}(t)$ are shown in blue. All eigenvalues decay to $\xi=0.5$ without crossing it. For both cases, we set the initial state to be the (trivial) ground state at $(J_{1}/J_{2},\gamma_{1}/(2J_{2}))=(1.5,0)$, the system size $L=20$, and the size of subsystem S is $M=10$ sites. Note crucially that, while the ES dynamics reflect the OBC spectrum, our physical system is under PBC.
@@
The model.— We investigate the dynamics of an open quantum system, illustrated in Fig. 1, where a free fermionic system ($\mathrm{S}+\bar{\mathrm{S}}$) is linearly coupled to a Markovian environment (${\rm env}$). After integrating out the environmental degrees of freedom, the reduced density matrix $\hat{\rho}={\rm Tr}_{\rm env}\hat{\rho}_{\rm tot}$ is governed by a quadratic Gorini-Kossakowski-Sudarshan-Lindblad (GKSL) equation [25, 26], given by,
$\partial_{t}\hat{\rho}(t)$ $=-i[\hat{H},\hat{\rho}(t)]+\sum_{\alpha}\mathcal{D}[\hat{L}_{\alpha}]\hat{\rho}(t)+\sum_{\alpha}\mathcal{D}[\hat{G}_{\alpha}]\hat{\rho}(t)$
=^$=\hat{\mathcal{L}}\hat{\rho}(t),$ (1)
@@
where the superoperator $\hat{\mathcal{L}}$ is the Liouvillian and $\hat{H}=\sum_{i,j=1}^{L}H_{i,j}\hat{c}_{i}^{\dagger}\hat{c}_{j}$ is the Hamiltonian of the relevant system $(\textrm{S}+\bar{\textrm{S}})$. Here, $\hat{c}_{i}^{\dagger}$ and $\hat{c}_{i}$ are creation and annihilation operators of fermions at site $i$, respectively. The dissipation arising from the coupling of the system to the Markovian reservoir is described by $\mathcal{D}[\hat{M}]\hat{\rho}:=\hat{M}\hat{\rho}\hat{M}^{\dagger}-\frac{1}{2}\{\hat{M}^{\dagger}\hat{M},\hat{\rho}(t)\}$. Each coupling to the Markovian reservoir, labeled by $\alpha$ (see Fig. 2(a)), induces incoherent addition or removal of particles and is described by linear jump operators, $\hat{L}_{\alpha}=\sum_{i}\ell_{i}^{\alpha}\hat{c}_{i}$ and $\hat{G}_{\alpha}=\sum_{i}g^{*\alpha}_{i}\hat{c}^{\dagger}_{i}$, respectively, where $\ell_{i}^{\alpha}$ and $g^{\alpha}_{i}$ are complex numbers.
@@
ES crossings as a non-Hermitian bulk-boundary correspondence indicator.— To dynamically detect the non-Hermitian bulk-boundary correspondence, we propose examining the ES dynamics after a dissipative quench. As depicted in Fig. 1, we partition the relevant system $(\mathrm{S}+\bar{\mathrm{S}})$ further into the subsystem ${\mathrm{S}}=\{1,2,...,M\}$ and its complement $\bar{\mathrm{S}}=\{M+1,M+2,...,L\}$. The reduced density matrix of the subsystem can be written as $\hat{\rho}_{\mathrm{S}}={\rm Tr}_{\bar{\mathrm{S}}}\hat{\rho}=(1/Z_{\mathrm{S}})e^{-\hat{H}_{\rm S}}$ with $Z_{\mathrm{S}}={\rm Tr}_{\mathrm{S}}[e^{-\hat{H}_{\rm S}}]$, where ${\rm Tr}_{\mathrm{S}(\bar{\mathrm{S}})}$ denotes the trace over subsystem S(${\mathrm{S}}(\bar{\mathrm{S}})$. We refer to the spectrum of $\hat{H}_{\rm S}$ as the ES $\epsilon_{l}$, the central quantity of interest. Using the property that $\hat{\rho}_{\rm S}$ is Gaussian, the entanglement Hamiltonian can be decomposed as $\hat{H}_{\rm S}=\sum_{l}\epsilon_{l}\hat{f}_{l}^{\dagger}\hat{f}_{l}$, where $\hat{f}_{l}$ is a fermionic annihilation operator of the $l$-th eigenmode. In the absence of the system-environment coupling, this recovers the conventional definition of the ES for closed fermionic systems, which is known to successfully detect bulk-boundary correspondence [27, 28, 29].
@@
References [30, 31, 32] considered the situation where they first prepared a state corresponding to the ground state of a Hamiltonian in the trivial phase followed by a sudden change in the parameters. They showed that when the post-quench Hamiltonian has a topologically non-trivial (trivial) ground state, (no) zero-crossings of the ES occur.
@@
We investigate below a similar setup where we also quench the system from a topologically trivial to a non-trivial or trivial state in an open quantum system governed by Eq. (1[4, 35, 36]. Importantly, the topology of our system is characterized by the non-Hermitian effective Hamiltonian $\hat{H}_{\rm eff}$, given by Eq. (2), that exhibits NHSE. We find that an analogous phenomenon to the closed-system counterpart occurs, which surprisingly reflects the topology of an OBC system even when the system itself is in PBC: physically traversing the system is unnecessary to detect topological phases of $\hat{H}_{\rm eff}$ defined under open boundary conditions.
@@
Practically, in our quadratic system, the ES can be computed from the covariance matrix $\mathsf{C}(t)$, whose $(j,i)$ component is given by the spatial correlation function $[{\sf C}]_{j,i}(t):={\rm Tr}_{\mathrm{S}+\bar{\mathrm{S}}}[\hat{\rho}(t)\hat{c}^{\dagger}_{i}\hat{c}_{j}]$ [40]. By taking advantage of the property that the density matrix is always Gaussian (as long as we take the initial state to be Gaussian) and Wick’s theorem applies as a result, the ES can be derived from the spectrum of the covariance matrix ${\sf C}_{\sf SS}$ in the subsystem S defined as the correlation matrix $\sf C$ whose indices are restricted to subsystem S,
$[{\sf C}_{\rm SS}]_{j,i}(t):={\rm Tr}_{\mathrm{S}+\bar{\mathrm{S}}}[\hat{\rho}(t)\hat{c}^{\dagger}_{i}\hat{c}_{j}];\;(i,j)\in{\sf S},$ (3)
@@
which can be expressed using the reduced density matrix of the subsystem S as $[{\sf C}_{\rm SS}]_{j,i}(t)={\rm Tr}_{\rm S}[\hat{\rho}_{\rm S}(t)\hat{c}^{\dagger}_{i}\hat{c}_{j}]$. Letting $\xi_{l}(t)$ be the eigenspectrum of $\mathsf{C}_{\rm SS}(t)$, one finds the relation between $\xi_{l}(t)$ and the ES $\epsilon_{l}(t)$  [40] (See Supplemental Materials (SM) [39]):
@@
Lindbladian SSH model.— To be specific, we consider the following system exhibiting LSE [5, 11] (Fig. 2(a)). For the system Hamiltonian, we consider the tight-binding Hamiltonian
$\hat{H}=\sum\limits_{j}\left(J_{1}\hat{c}_{A,j}^{\dagger}\hat{c}_{B,j}+J_{2}\hat{c}^{\dagger}_{B,j}\hat{c}_{A,j+1}+J_{3}\hat{c}_{A,j}^{\dagger}\hat{c}_{B,j+1}\right)+h.c.$ (6)
For $J_{3}=0$, the Hamiltonian reduces to the SSH model [37]. The dissipation of our model is given by the following jump operators for $j=1,2,\cdots,N$ (see Fig. 2 (a)): $\hat{L}_{j,1}=\sqrt{\frac{\gamma_{1}}{2}}\left(\hat{c}_{A,j}+i\hat{c}_{B,j}\right)$, $\hat{G}_{j,1}=\sqrt{\frac{\gamma_{1}}{2}}(\hat{c}_{A,j}^{\dagger}-i\hat{c}_{B,j}^{\dagger})$, $\hat{L}_{j,2}=\sqrt{\frac{\gamma_{2}}{2}}\left(\hat{c}_{B,j}+i\hat{c}_{A,j+1}\right)$, $\hat{G}_{j,2}=\sqrt{\frac{\gamma_{2}}{2}}(\hat{c}_{B,j}^{\dagger}-i\hat{c}_{A,j+1}^{\dagger})$. Here, we have chosen jump operators such that $\hat{L}_{j,s=1,2}^{\dagger}=\hat{G}_{j,s=1,2}$, ensuring the system reaches an infinite-temperature state at the long-time limit. The ES asymptotically converges to zero $\epsilon_{n}=0$ $(\xi_{n}=0.5)$, avoiding the ES from crossing zero for trivial reasons with no topological origin [4, 5].
This choice of Hamiltonian and jump operators yield the effective Hamiltonian (Eq. (2)),
@@
$+J_{3}\hat{c}_{A,j}^{\dagger}\hat{c}_{B,j+1}\Big{)}+h.c.\Big{]}+\sum_{j}\Big{(}\frac{\gamma_{1}}{2}(\hat{c}_{A,j}^{\dagger}\hat{c}_{B,j}-\hat{c}^{\dagger}_{B,j}\hat{c}_{A,j})$
$+\frac{\gamma_{2}}{2}(\hat{c}^{\dagger}_{B,j}\hat{c}_{A,j+1}-\hat{c}_{A,j+1}^{\dagger}\hat{c}_{B,j})\Big{)}-i\frac{\gamma_{1}+\gamma_{2}}{2}\hat{n},$ (7)
with $\hat{n}:=\sum_{j}[\hat{c}^{\dagger}_{A,j}\hat{c}_{A,j}+\hat{c}^{\dagger}_{B,j}\hat{c}_{B,j}]$, which is the non-Hermitian SSH model well studied in the literature (except for the constant shift in the imaginary axis that does not affect its topological features) which is known to exhibit NHSE [5, 6]. For PBC, the Hamiltonian $\hat{H}_{\rm eff}$ can expressed as $\hat{H}_{\rm eff}=\sum_{k}\bm{c}^{\dagger}(k)\mathsf{H}_{\rm eff}(k)\bm{c}(k)$, where $\bm{c}(k):=(\hat{c}_{A}(k),\hat{c}_{B}(k))^{T}$ with $\hat{c}_{A(B)}(k)=1/\sqrt{L}\sum_{j}e^{-ikj}\hat{c}_{A(B),j}$ and $\mathsf{H}_{\rm eff}(k)=h_{1}(k)\sigma_{x}+h_{2}(k)\sigma_{y}$, where $h_{1}(k)=J_{1}+(J_{2}+J_{3})\cos(k)-i\gamma_{2}/2\sin(k)$, $h_{2}(k)=(J_{2}-J_{3})\sin(k)+i\gamma_{2}/2\cos(k)-i\gamma_{1}/2$. Here, $\sigma_{i}$ with $i\in\{x,y,z\}$ are the Pauli matrices and we have ignored the constant shift term. Since the effective Hamiltonian has a sublattice symmetry $\sigma_{z}\mathsf{H}_{\rm eff}(k)\sigma_{z}^{-1}=-\mathsf{H}_{\rm eff}(k)$, we can define the ${\mathbb{Z}}$ topological number of $\mathsf{H}_{\rm eff}(k)$, given by the winding number $\omega_{\rm PBC}=\oint\frac{dk}{4\pi i}{\rm Tr}[\sigma_{z}\mathsf{H}_{\rm eff}^{-1}(k)\frac{d}{dk}\mathsf{H}_{\rm eff}(k)]$ [2]. The winding number can be similarly defined for OBC, $\omega_{\rm OBC}$, by generalizing the momentum integral in the PBC winding number, $\omega_{\rm PBC}$, to those with complex values along the generalized Brillouin zone. (See e.g., Refs. [6, 20] for details.)
@@
In our simulation, we set the pre-quench state to be the ground state in the trivial phase of Eq. (7), in the Hermitian limit $\gamma_{1}=\gamma_{2}=0$ with $J_{3}=0$.